Consider two​ economies, A and B. Economy A has a marginal propensity to consume of​ 0.9, a net tax rate of 0.1 and a marginal propensity to import of 0.1. Economy B has a marginal propensity to consume of​ 0.6, a net tax rate of 0.2 and a marginal propensity to import of 0.2. Suppose there is a decrease in autonomous investment of​ $5 billion in each of these economies. Which of the following statements is​ true?

A.The AD curve shifts to the right the same amount in both economies.

B.The AD curve shifts farther to the right in Economy A than Economy B.

C.The AD curve shifts to the left the same amount in both economies.

D.The AD curve shifts farther to the left in Economy A than Economy B.

E.The simple multiplier is larger in Economy B.
Answer #1

The formula for Fiscal multiplier for investment is 1/(1-MPC(1-T)+MPI) where MPC=marginal propensity to consume, MPI=marginal propensity to import, T=tax rate

For A, the multiplier is 1/(1-0.9(1-0.1)+0.1)=1/(1.1-0.81)=1/0.29=3.45

For B, the multiplier is 1/(1-0.6(1-0.2)+0.2)=1/(1.2-0.48)=1/0.72=1.39

Since the multiplier is higher for A, AD curve will shift farther for A than B and the shift will be towards right as there is new investment.

So, the correct answer is B.